OpenStudy (anonymous):

Evaluate 3^2 + (8 - 2) • 4 - 6 over 3

OpenStudy (anonymous):

GEMDAS: Grouping: 8-2=6. Exponents: N/A Multiplication: 6*4=24. Division: Saved for last because everything else is technically within the group of the numerator. Addition: 32+24=56 Subtraction: 56-6=50. Division: 50/3, cannot be rationally simplified further. Unless the 'over 3' only applies to the 6.

OpenStudy (anonymous):

Wait, did 32 just change to 3^2, or am I crazy?

OpenStudy (anonymous):

I just changed it, and the 'over 3' only applies to the 6.

OpenStudy (anonymous):

Well, that changes things significantly.

OpenStudy (anonymous):

I'm sorry. >.<

OpenStudy (anonymous):

Grouping: 8-2=6. Exponents: 3^2=9. Multiplication: 6*4=24 Division: 6/3=2. Addition and Subtraction: 9 + 24 - 2=31.

OpenStudy (anonymous):

Just make sure to apply parentheses next time, otherwise there will be a slew of different possible answers.

OpenStudy (anonymous):

Dude, thank you so much. I appreciate it. And can you show me an example of how I should of asked my question correctly?

OpenStudy (anonymous):

(3^2)+(8-2)*4-(6/3) Parentheses around the exponent the number is raised to, so that you don't think there's a larger value in the exponent of three. Put the 6/3 in parenthesis or you'll think more of it is being divided by three. Further, you should keep the way you type your operations consistent: don't type partially in symbols and then for division type 'over three', because that's confusing. I hope this helps, it's a LOT easier to answer these questions when you use parentheses right.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem, have a nice day.